# Distillation Learning Guided by Image Reconstruction for One-Shot Medical Image Segmentation

### v1

@@ 1 Introduction
2 Related Work
2.1 One-Shot Medical Image Segmentation
2.1.1 Atlas-Based Segmentation (ABS)
@@
2.1.2 Learning Registration to Learn Segmentation (LRLS)
@@ 3 Methods
3.1 Data Augmentation with Image Registration and Contrastive Learning
@@
3.1.1 Registration-based Data Augmentation Network
In this work, we use a variant of CLMorph. Unlike the original model (CLMorph [1]), we modified the encoder network to eliminate the reparameterization trick for Gaussian distribution, simplifying the model by directly encoding the input image without introducing random variability in the latent space. We doubled the number of convolution filters to enhance feature extraction efficiency. Additionally, we expanded the original model’s similarity loss and smooth loss to better handle the intensity and deformation variability across datasets of different organs and modalities (see Section 3.3.1 for more details). Note that the registration network is not the focus of this work, as our framework adapts to any unsupervised registration network.
@@ 3.1.1 Registration-based Data Augmentation Network
3.1.2 Feature-Level Contrastive Learning
@@
3.2 Feature Distillation Learning with Image Reconstruction for Segmentation
@@
3.2.1 Teacher-Student Network
@@
3.2.2 Feature Distillation Learning
@@
Previous studies indicate that features closer to the output are highly correlated with the output task [45]. Based on this, we experimentally analyzes the number of feature layers used for distillation learning, calculating hint loss for features from the last 1-5 layers of both teacher and student networks. As detailed in Section 4.5.3, we found that selecting the last two layers yielded the best Dice coefficient for the results reported in this work. In Section 4.5.2, we further studied the impact of different distillation losses on our framework. Instead of using the commonly used L2 norm [34], we employ cosine similarity to decrease the feature distance, thereby enhancing the student network’s performance with the following loss:
@@
3.3 Optimization and Inference Stage
3.3.1 Optimization of Registration-based Data Augmentation
@@
3.3.2 Optimization of Feature Distillation Learning
For the teacher network, we compute the mean squared error (MSE) similarity between the real image and its reconstructed counterpart as the reconstruction loss, aiming to enhance the overall reconstruction performance, as expressed in the following:
@@ 3.3.3 Inference
4 Experiments and Results
Table 1: Quantitative evaluation on the OASIS, BCV, and VerSe datasets used the Dice Similarity Coefficient (DSC) and the 95th percentile Hausdorff Distance (HD95mm) metrics. The best result, excluding the upper limit, is highlighted in blue.
@@ 4.1 Datasets and Preprocessing
3) Large Scale Vertebrae Segmentation Challenge (VerSe): a subset of the MICCAI VerSe20 Challenge dataset [55], containing 79 thoracic vertebrae CT images: 1 atlas, 69 training, and 16 test images with labels for 13 anatomical structures. Pre-processing with ANTs [22] includes rigid pre-alignment, cropping, padding, resampling, and windowing, resulting in a resolution of 64 × 96 × 192 with $2mm$ isotropic voxels.
@@ 4.2 Experimental Settings
Figure 5: Comparison between our method and SOTA one-shot MIS methods for an arbitrary case in each dataset. Yellow boxes highlight regions where our method outperforms the SOTA methods.
4.3 Implementation Details and Evaluation Metrics
@@
4.4 Comparisons with Leading Methods
In Table 1, we compare our method with 12 others (Section 4.2) on the OASIS, BCV, and VerSe datasets using DSC and HD95mm metrics (Section 4.3). Three key observations are: 1) LS methods show adaptability to one-shot segmentation but have poor segmentation quality (high HD95) and unstable performance across labels. They perform well on large organs like the cerebral white matter but fail on smaller structures like vessels. 2) ABS methods perform stably for small structures due to consistent anatomical features. For OASIS dataset, they achieve nearly 80% DSC, only 5.4% lower than ours. However, their performance is limited by atlas-test image similarity and deformation extent, struggling with large deformations in BCV and VerSe tasks. Notably, the traditional ABS method SyN outperforms deep learning ABS methods in these tasks. 3) LRLS methods significantly outperform LS and ABS methods, thanks to deep learning and extensive training data, achieving over 80% DSC across all tasks.
Table 2: Comparison of segmentation performance for 35 brain structures on the OASIS dataset. Fully supervised serves as the upper bound benchmark. WM: white matter, CX: cortex, Vent: ventricle, STD: standard deviations. Scores for symmetrical brain regions are combined. The best result, excluding the upper limit, is highlighted in blue.
@@ 4.4 Comparisons with Leading Methods
Compared to other LRLS-based models, our method excels in two key areas. Firstly, it achieves the highest DSC and lowest HD95 across all three datasets, demonstrating excellent performance and adaptability across different modalities and organs. Secondly, guided by distillation learning, it significantly enhances its understanding of anatomical features, leading to good segmentation performance for most organs. Segmentation details for each label in the OASIS (Table 2) and BCV (Fig. 4) datasets showcase the superior anatomical feature understanding of our framework.
@@ 4.5 Ablation Studies
4.5.1 Impact of Model Architecture Components
Our model has three key components, as outlined in Table 3. The ABS method (baseline) achieves a DSC of 79.9% and an HD95 of 3.214 mm. Adding the LRLS concept and training the student network with synthetic data (M2) slightly improves DSC by 0.7%, but HD95 increases by 0.611 mm. This limited improvement is due to the lack of teacher network guidance. With distillation learning (Ours), DSC improves with a 4.8% and HD95 decreases by 2.109 mm.
Table 3: Ablation study of model architecture components. SI+SL trains the student network with synthetic images and labels, while RI+SL uses real images with synthetic labels.
@@ 4.5.1 Impact of Model Architecture Components
Additionally, Table 3 highlights the necessity of using synthetic images and labels (SI+SL) aligned via atlas distortion, enhancing learning efficiency. Conversely, using real images with synthetic labels (RI+SL) can introduce inaccuracies due to potential misalignment [2]. This is evident in M3, where RI+SL results in a 3.7% decrease in DSC. Exposing the student network to real images does not necessarily improve performance. In M1, learning from RI+SL results in a DSC 1.8% lower than M2 (SI+SL) and even lower than the baseline, underscoring the importance of alignment and the teacher network. Overall, the carefully designed components of our method are crucial for achieving the best performance.
4.5.2 Impact of Integrated Distillation Learning
Our feature distillation learning process (Eq. (12)) includes the segmentation loss $\mathcal{L}_{seg}$, the hint loss $\mathcal{L}_{hint}$, and the reconstruction loss $\mathcal{L}_{recon}$. Table 4 demonstrates the effectiveness of each component. The results show that incorporating hint loss and reconstruction loss significantly enhances the student network’s segmentation ability, indicating that our mix of loss functions ensures optimal segmentation results.
Table 4: Impact of integrated distillation learning module.
@@ 4.5.2 Impact of Integrated Distillation Learning
To further assess the significance of our proposed hint loss (Eq. (3)), we conducted an ablation study to examine the segmentation performance using $\mathcal{L}_{hint}$ with L2 norm and cosine similarity. As shown in Table 5, employing cosine similarity as the hint loss outperforms L2 loss within the distillation learning framework. These findings confirm that using cosine similarity reduces the feature distance, thereby enhancing the segmentation performance of the student network. Fig. 6 shows the necessity of our designed teacher-student network by comparing the feature maps for all three datasets. For accurate comparisons, all compared models share the same network architecture.
Table 5: Comparison of different hint losses in distillation learning.
@@
Figure 6: Feature maps from the last layer of each network for an arbitrary case in the OASIS, BCV, and VerSe datasets. ResUNet is the student network without teacher guidance. Red boxes highlight feature maps of the student network that closely resemble the teacher’s.
4.5.3 Impact of Hyperparameters
@@
5 Discussion
ABS methods [24, 25, 14, 26, 15, 1] generate predicted labels by propagating atlas labels through registration. However, their segmentation accuracy is unstable due to reliance on the similarity between atlas and target images, and they lack robustness for segmenting large organs with significant deformations, such as in abdominal CT (Figs. 4-5). These findings are consistent with previous literature [28].
In our experiments, LRLS methods [2, 3] consistently outperformed ABS methods in segmentation tasks. Current LRLS methods either synthesize richer pseudo-datasets through carefully designed sampling strategies [2, 30, 3] or enhance performance under small sample conditions via joint registration-segmentation optimization [3, 29, 31]. Despite their effectiveness, these approaches often limit the quality and quantity of synthetic images when faced with challenging or scarce unlabeled data. Joint optimization models are parameter-heavy, difficult to train, and have limited adaptability to different modalities and organs. Additionally, they fail to fully utilize real unlabeled images, which contain more reliable anatomical structures than synthetic images.
@@ 5 Discussion
While our method showed promising results, it has several limitations. Firstly, like all one-shot MIS methods, it relies on limited labeled samples, hindering the model’s ability to capture complex anatomical structures and diverse image features due to high medical imaging heterogeneity. This can cause significant discrepancies between test data from unknown domains and the training data, affecting accuracy and robustness. Additionally, the framework struggles with surgically altered anatomical structures, as it is trained on unaltered cases. Finally, further assessment of our method’s generalizability to more diverse pathological images is needed.
Future research should integrate multi-center, cross-domain datasets with multi-modal data to improve robustness and accuracy. Enhanced data augmentation, advanced network architectures, incorporating domain knowledge, and generalization techniques can further improve effectiveness and applicability.
6 Conclusion

### v2

@@ I Introduction
II Related Work
II-A One-Shot Medical Image Segmentation
II-A1 Atlas-Based Segmentation (ABS)
@@
II-A2 Learning Registration to Learn Segmentation (LRLS)
@@ III Methods
III-A Data Augmentation with Image Registration and Contrastive Learning
@@
III-A1 Registration-based Data Augmentation Network
In this work, we use a variant of CLMorph. Unlike the original model (CLMorph [1]), we modified the encoder network to eliminate the reparameterization trick for Gaussian distribution, simplifying the model by directly encoding the input image without introducing random variability in the latent space. We doubled the number of convolution filters to enhance feature extraction efficiency. Additionally, we expanded the original model’s similarity loss and smooth loss to better handle the intensity and deformation variability across datasets of different organs and modalities (see Section III-C1 for more details). Note that the registration network is not the focus of this work, as our framework adapts to any unsupervised registration network.
@@ III-A1 Registration-based Data Augmentation Network
III-A2 Feature-Level Contrastive Learning
@@
III-B Feature Distillation Learning with Image Reconstruction for Segmentation
@@
III-B1 Teacher-Student Network
@@
III-B2 Feature Distillation Learning
@@
Previous studies indicate that features closer to the output are highly correlated with the output task [45]. Based on this, we experimentally analyzes the number of feature layers used for distillation learning, calculating hint loss for features from the last 1-5 layers of both teacher and student networks. As detailed in Section IV-E3, we found that selecting the last two layers yielded the best Dice coefficient for the results reported in this work. In Section IV-E2, we further studied the impact of different distillation losses on our framework. Instead of using the commonly used L2 norm [34], we employ cosine similarity to decrease the feature distance, thereby enhancing the student network’s performance with the following loss:
@@
III-C Optimization and Inference Stage
III-C1 Optimization of Registration-based Data Augmentation
@@
III-C2 Optimization of Feature Distillation Learning
For the teacher network, we compute the mean squared error (MSE) similarity between the real image $y$ and its reconstructed counterpart $y^{r}$ as the reconstruction loss, aiming to enhance the overall reconstruction performance, as expressed in the following:
@@ III-C3 Inference
IV Experiments and Results
TABLE I: Quantitative evaluation on the OASIS, BCV, and VerSe datasets used the Dice Similarity Coefficient (DSC) and the 95th percentile Hausdorff Distance (HD95mm) metrics. The best result, excluding the upper limit, is highlighted in bold.
@@ IV-A Datasets and Preprocessing
3) Large Scale Vertebrae Segmentation Challenge (VerSe): a subset of the MICCAI VerSe20 Challenge dataset [55], containing 86 thoracic vertebrae CT images: 1 atlas, 69 training, and 16 test images with labels for 13 anatomical structures. Pre-processing with ANTs [22] includes rigid pre-alignment, cropping, padding, resampling, and windowing, resulting in a resolution of 64 × 96 × 192 with $2mm$ isotropic voxels.
@@ IV-B Experimental Settings
Figure 5: Comparison of our method with SOTA one-shot MIS methods on arbitrary cases. Yellow boxes indicate regions where our method is superior.
IV-C Implementation Details and Evaluation Metrics
@@
IV-D Comparisons with Leading Methods
In Table I, we compare our method with 12 others (Section IV-B) on the OASIS, BCV, and VerSe datasets using DSC and HD95mm metrics (Section IV-C). Three key observations are: 1) LS methods show adaptability to one-shot segmentation but have poor segmentation quality (high HD95) and unstable performance across labels. They perform well on large organs like the cerebral white matter but fail on smaller structures like vessels. 2) ABS methods perform stably for small structures due to consistent anatomical features. For OASIS dataset, they achieve nearly 80% DSC, only 5.4% lower than ours. However, their performance is limited by atlas-test image similarity and deformation extent, struggling with large deformations in BCV and VerSe tasks. Notably, the traditional ABS method SyN outperforms deep learning ABS methods in these tasks. 3) LRLS methods significantly outperform LS and ABS methods, thanks to deep learning and extensive training data, achieving over 80% DSC across all tasks.
TABLE II: Comparison of segmentation performance for 35 brain structures on the OASIS dataset. Fully supervised serves as the upper bound benchmark. WM: white matter, CX: cortex, Vent: ventricle, STD: standard deviations. Scores for symmetrical brain regions are combined. The best result, excluding the upper limit, is highlighted in bold.
@@ IV-D Comparisons with Leading Methods
Compared to other LRLS-based models, our method excels in two key areas. Firstly, it achieves the highest DSC and lowest HD95 across all three datasets, demonstrating excellent performance and adaptability across different modalities and organs. Secondly, guided by distillation learning, it significantly enhances its understanding of anatomical features, leading to good segmentation performance for most organs. Segmentation details for each label in the OASIS (Table II) and BCV (Fig. 4) datasets showcase the superior anatomical feature understanding of our framework.
@@ IV-E Ablation Studies
IV-E1 Impact of Model Architecture Components
Our model has three key components, as outlined in Table III. The ABS method (baseline) achieves a DSC of 79.9% and an HD95 of 3.214 mm. Adding the LRLS concept and training the student network with synthetic data (M2) slightly improves DSC by 0.7%, but HD95 increases by 0.611 mm. This limited improvement is due to the lack of teacher network guidance. With distillation learning (Ours), DSC improves with a 4.8% and HD95 decreases by 2.109 mm.
TABLE III: Ablation study of model architecture components. SI+SL trains the student network with synthetic images and labels, while RI+SL uses real images with synthetic labels.
@@ IV-E1 Impact of Model Architecture Components
Additionally, Table III highlights the necessity of using synthetic images and labels (SI+SL) aligned via atlas distortion, enhancing learning efficiency. Conversely, using real images with synthetic labels (RI+SL) can introduce inaccuracies due to potential misalignment [2]. This is evident in M3, where RI+SL results in a 3.7% decrease in DSC. Exposing the student network to real images does not necessarily improve performance. In M1, learning from RI+SL results in a DSC 1.8% lower than M2 (SI+SL) and even lower than the baseline, underscoring the importance of alignment and the teacher network. Overall, the carefully designed components of our method are crucial for achieving the best performance.
IV-E2 Impact of Integrated Distillation Learning
Our feature distillation learning process (Eq. (12)) includes the segmentation loss $\mathcal{L}_{seg}$, the hint loss $\mathcal{L}_{hint}$, and the reconstruction loss $\mathcal{L}_{recon}$. Table IV demonstrates the effectiveness of each component. The results show that incorporating hint loss and reconstruction loss significantly enhances the student network’s segmentation ability, indicating that our mix of loss functions ensures optimal segmentation results.
TABLE IV: Impact of integrated distillation learning module.
@@ IV-E2 Impact of Integrated Distillation Learning
To further assess the significance of our proposed hint loss (Eq. (3)), we conducted an ablation study to examine the segmentation performance using $\mathcal{L}_{hint}$ with L2 norm and cosine similarity. As shown in Table V, employing cosine similarity as the hint loss outperforms L2 loss within the distillation learning framework. These findings confirm that using cosine similarity reduces the feature distance, thereby enhancing the segmentation performance of the student network. Fig. 6 shows the necessity of our designed teacher-student network by comparing the feature maps for all three datasets. For accurate comparisons, all compared models share the same network architecture.
TABLE V: Comparison of different hint losses in distillation learning.
@@
Figure 6: Feature maps from the last layer of each network for three randomly selected channels of an arbitrary sample in the OASIS, BCV, and VerSe datasets. ResUNet is the student network without teacher guidance. Red boxes highlight feature maps of the student network that closely resemble the teacher’s.
IV-E3 Impact of Hyperparameters
@@
V Discussion
ABS methods [24, 25, 14, 26, 15, 1] generate predicted labels by propagating atlas labels through registration. Their segmentation accuracy tends to unstable due to reliance on the similarity between atlas and target images, and they lack robustness for segmenting large organs with significant deformations, as shown in abdominal CT (Figs. 4-5). These findings are consistent with previous literature [28].
In our experiments, the LRLS methods [2, 3] consistently outperformed ABS methods in segmentation tasks. Current LRLS methods either synthesize richer pseudo-datasets through carefully designed sampling strategies [2, 30, 3] or enhance performance under small sample conditions via joint registration-segmentation optimization [3, 29, 31]. Despite their effectiveness, these approaches often limit the quality and quantity of synthetic images when faced with challenging or scarce unlabeled data. Joint optimization models are parameter-heavy, difficult to train, and have limited adaptability to different modalities and organs. Additionally, they fail to fully utilize real unlabeled images, which contain more reliable anatomical structures than synthetic images.
@@ V Discussion
While our method demonstrated promising results, it has several limitations. Medical images are highly heterogeneous, and labeling anatomical structures is both costly and time-intensive. As with other one-shot MIS methods, our framework is currently tailored for scenarios with limited labeled samples. This constraint may hinder the model’s ability to capture complex anatomical structures and diverse imaging features, potentially leading to performance discrepancies between training data and unseen test data from unknown domains. Future work will address these challenges by evaluating the generalizability of our method on larger datasets with more comprehensive annotations. For instance, the proposed method could be extended to segment various organs (e.g., brain, abdomen and vertebrae) for diverse clinical applications. Specifically, we are actively working on applying the framework to a large dataset of pediatric CT images to segment whole brain structures, aiming to support the assessment of brain diseases and developmental processes.
VI Conclusion